OpenStudy (anonymous):

If the force acting on an object stays the same, then the acceleration of the object is inversely proportional to its mass. If an object with a mass of 9 kilograms accelerates at a rate of 6 meters per second per second by a force, find the rate of acceleration of an object with a mass of 2 kilograms that is pulled by the same force. @zzr0ck3r @Mehek14 @Conqueror

OpenStudy (anonymous):

F=ma m1a1=m2a2

OpenStudy (anonymous):

What do the variables mean?

OpenStudy (anonymous):

@robtobey @mathmate can you guys help me please?

OpenStudy (anonymous):

m1=mass a1=acceleration similarly for m2 and a2

OpenStudy (anonymous):

,m1,a1 for first body and m2,a2 for second body.

OpenStudy (anonymous):

Okay so m1a1= 9 & 6

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

m2= 2?

OpenStudy (anonymous):

yes ,find a2

OpenStudy (anonymous):

How do i find a2?

OpenStudy (anonymous):

9*6=2*a2

OpenStudy (anonymous):

(9)(6)=(2)(a) 54=2a correct?

OpenStudy (anonymous):

yes a=?

OpenStudy (anonymous):

then i divide by 2.. 54/2=2a/2 correct?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

a=27

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

But my choices are... A.5 inches B. 4 inches C.4.5 inches D. 5.5 inches

OpenStudy (anonymous):

check your choices in the question kg and meters how can be inches?

OpenStudy (anonymous):

Opps wrong question sorry...

OpenStudy (anonymous):

The correct answer is B. 27 meters per second. can you help me with one more please?

OpenStudy (anonymous):

A rain gutter is made from sheets of aluminum that are 18 inches wide by turning up the edges to form right angles. Determine the depth of the gutter that will maximize its cross-sectional area and allow the greatest amount of water to flow. A. 5 inches B. 4 inches C. 4.5 inches D. 5.5 inches

OpenStudy (anonymous):

i am leaving now for sometime.

OpenStudy (anonymous):

Okay bye thanks.